VIBRATIONAL PREDISSOCIATION OF Na^{+}(CH_{3}OH)_{m}((CH_{3})_{2}CO)_{n} (m + n > 1; m, n ≥ 1) FROM 2.5 TO 3.0 MICRONS.

\underline{C. J. WEINHEIMER} AND J. M. LISY, Department of Chemistry, University of Illinois at Urbana-Champaign,Urbana, IL 61801.

The vibrational predissociation spectra of Na^{+}(CH_{3}OH)_{m}((CH_{3})_{2}CO)_{n} (m + n > 1; m, n ≥ 1), recorded between 3250 - 3800 cm^{−1}, will be discussed and compared with clusters, Na^{+}(CH_{3}OH)_{m}, that do not contain any acetone molecules (n=0). Acetone is not infrared active in this spectral region, but is capable of participating in strong hydrogen bonds with methanol. Any changes in the spectra between Na^{+}(CH_{3}OH)_{m} and Na^{+}(CH_{3}OH)_{m}((CH_{3})_{2}CO)_{n} can be attributed to the influence of one or more acetone molecules. As in the case of Na^{+}(CH_{3}OH)_{m} clusters, the smaller cluster sizes show an absorption around 3670 cm^{−1}, which is attributed to methanols with ‘free’ O-H groups. Methanols acting as proton donors to other acetone or methanol molecules absorb below 3500 cm^{−1}. The structural impact due to the presence of one or more acetones will be discussed.